OpenStudy (nana82058):

Ms. Chang has 24 students in her class, and 6 of them are boys. What is the ratio of boys to girls?

OpenStudy (aaronandyson):

This one is really simple.

OpenStudy (aaronandyson):

Total no of students = 24 No of boys = 6 No of girls = 24-6 = 18 Ratio 18:6 3:1

OpenStudy (mathstudent55):

No and no.

OpenStudy (mathstudent55):

Boys: 6 Total: 24 Girls: 24 - 6 = 18 Boys to girls = 6:18 = 1:3

OpenStudy (mathstudent55):

Both of you know how to express a ratio and to reduce it, but you must be careful when you read a problem to make sure you answer the question that was asked.

OpenStudy (nana82058):

thanks @mathstudent55 this help a lot

OpenStudy (mathstudent55):

You're welcome.